# Segmenting Action-Value Functions Over Time-Scales in SARSA using TD($\Delta$)

### v1

@@ 5.2 Evaluation for Minimizing $k_{z}$ Values in SARSA via Phased Updates
$\Delta_{t}^{\hat{Q}_{\gamma}}\leq\epsilon\bigg{(}\underbrace{\frac{1-\gamma^{k}}{1-\gamma}\bigg{)}}_{\text{variance term}}+\underbrace{\gamma^{k}\Delta_{t-1}^{\hat{Q}_{\gamma}}}_{\text{bias term}}$ (25)
@@
In Eq. 25, A variance term ($\frac{1-\gamma^{k}}{1-\gamma}$) is arising due to sampling error from rewards collected along trajectories. In particular, $\epsilon$ bounds the deviation of the empirical average of rewards from the true expected reward. Second term ($\gamma^{k}\Delta_{t-1}^{\hat{Q}_{\gamma}}$) in Eq. 25 is a bias term that arises from bootstrapping off the previous phase’s estimates. Similarly, we consider a phased SARSA version of multi-step $TD(\Delta)$. For each phase t, we update each W as follows:
$\hat{W}_{z,t}(s,a)=\frac{1}{n}\sum_{j=1}^{n}\bigg{(}\sum_{i=1}^{k-1}(\gamma_{z}^{i}-\gamma_{z-1}^{i})r_{i}^{(j)}+(\gamma_{z}^{k_{z}}-\gamma_{z-1}^{k_{z}})Q_{\gamma_{z-1}}(s_{t+k}^{(j)},a_{t+k}^{(j)})+\gamma_{z}^{k_{z}}\hat{W}_{z}(s_{t+k}^{(j)},a_{t+k}^{(j)})$ (26)
We now determine the phased error’s upper bound. The sum of the errors made by each W component $\sum_{z=0}^{Z}\Delta_{t}^{z}$ is known as $TD(\Delta)$, where $\Delta_{t}^{z}=max_{s,a}\bigg{\{}\bigg{|}\hat{W}_{z}(s,a)-W_{z}(s,a)\bigg{|}\bigg{\}}$. Key distinctions between SARSA and Q-learning in terms of the error bound: In SARSA, the error is determined by the agent’s actual actions, which means that the Q-values are updated according to the state-action pairs that the agent visits during exploration. The error in Q-learning is derived from the greedy action, which is the action that maximizes the Q-value for the next state, under the assumption that the agent always acts optimally.
@@
Assume that $\gamma_{0}\leq\gamma_{1}\leq\gamma_{2}\leq.....\leq\gamma_{z}=\gamma$ and $K_{0}\leq K_{1}\leq...\leq K_{z}=K$, for any $10<\delta<$ 1, let $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$, with probability $1-\delta$,
$\sum_{z=0}^{Z}\Delta_{t}^{z}\leq\epsilon\bigg{(}\frac{1-\gamma^{k}}{1-\gamma}\bigg{)}+\epsilon\bigg{(}\underbrace{\sum_{z=0}^{Z-1}\frac{\gamma_{z}^{k_{z+1}}-\gamma^{k_{z}}_{z}}{1-\gamma_{z}}\bigg{)}}_{\text{variance reduction}}+\underbrace{\sum_{z=0}^{Z-1}\bigg{(}\gamma^{k_{z}}_{z}-\gamma^{k_{z+1}}_{z}\bigg{)}\sum_{u=0}^{z}\Delta_{t-1}^{u}}_{\text{bias introduction}}+\gamma^{k}\sum_{z=0}^{Z}\Delta_{t-1}^{z}$ (27)
@@
The author examines the comparison of the limits for phased TD($\lambda$) in theorem 3 with those for phased TD($\Delta$) in theorem 4. This comparison demonstrates that phased TD($\Delta$) facilitates variance reduction equivalent to $\epsilon\sum_{z=0}^{Z-1}\frac{\gamma_{z}^{k_{z+1}}-\gamma^{k_{z}}_{z}}{1-\gamma_{z}}\leq 0$ but introduces a potential bias quantified as $\sum_{z=0}^{Z-1}\bigg{(}\gamma^{k_{z}}_{z}-\gamma^{k_{z+1}}_{z}\bigg{)}\sum_{u=0}^{z}\Delta_{t-1}^{u}\geq 0$. Utilizing phased TD($\Delta$) to decrease $k_{z}$ synchronizes updates more effectively with recent actions, hence lowering variance from high-discounted return components. Shortened $k_{z}$ values accrue bias more swiftly, as SARSA bootstraps from current policy estimations that may vary over phases. It is observed that when all $k_{z}$ values are identical, both algorithms yield the same upper bound.
@@
In SARSA, the return estimate encompasses subsequent states and actions dictated by the policy. The expected discounted return over T steps closely approximates the infinite-horizon discounted expected return after $\mathrm{T}$, where $\mathrm{T}\approx\frac{1}{1-\gamma}$[31]. Consequently, we can effectively simplify $k_{z}$ for any $\gamma_{z}$ such that $k_{z}\approx\frac{1}{1-\gamma_{z}}$, thereby adhering to this principle. Therefore, with $\mathrm{T}$ samples, similar to TD($\Delta$)[8], SARSA employing TD($\Delta$) utilizes $k_{z}\approx\frac{1}{1-\gamma_{z}}$ to establish a suitable balance between bias and variance over all time scales significantly less than $\mathrm{T}$. By establishing $k_{z}$ for each $\gamma_{z}$, where $\gamma_{z}^{\frac{1}{1-\gamma_{z}}}\leq\frac{1}{e}$, the approach guarantees that the effective horizon and variance are maintained within limits without requiring extensive parameter optimization. Each increase in $W_{z}$ facilitates the doubling of the effective horizon, hence ensuring logarithmic scaling with the quantity of action-value functions. SARSA can utilize this to modify additional parameters in alignment with the specified time scales, enhancing the algorithm’s efficiency in calculating long-term rewards without added complexity.
6 Experiments
We illustrate the efficacy of SARSA($\Delta$) via tests on multiple RL benchmarks, including the Ring MDP[9] and dense reward Atari environments. The findings indicate that SARSA($\Delta$) attains more rapid convergence and superior long-term efficacy relative to conventional SARSA and TD learning. All hyperparameter settings, extended details, and the reproducibility checklist for machine learning research[32] are included in the supplemental materials of this paper[8].
@@ 7 Conclusion
SARSA($\Delta$) expands SARSA by including multiple discount factors through TD($\Delta$), resulting in improved performance and stability. This approach for a more precise breakdown of future rewards while maintaining the benefits of on-policy learning. The Q-Learning version of TD($\Delta$) is a simple extension of this. We put it off for later.

### v2

@@ 5.2 Evaluation for Minimizing $k_{m}$ Values in SARSA via Phased Updates
$\Delta_{n}^{\hat{Q}_{\eta}}\leq\epsilon\bigg{(}\underbrace{\frac{1-\eta^{k}}{1-\eta}\bigg{)}}_{\text{variance term}}+\underbrace{\eta^{k}\Delta_{n-1}^{\hat{Q}_{\eta}}}_{\text{bias term}}$ (28)
@@
In Eq. 28, A variance term ($\frac{1-\eta^{k}}{1-\eta}$) is arising due to sampling error from rewards collected along trajectories. In particular, $\epsilon$ bounds the deviation of the empirical average of rewards from the true expected reward. Second term ($\eta^{k}\Delta_{n-1}^{\hat{Q}_{\eta}}$) in Eq. 28 is a bias term that arises from bootstrapping off the previous phase’s estimates. Similarly, we consider a phased SARSA version of multi-step $TD(\Delta)$. For each phase n, we update each D as follows:
$\hat{D}_{m,n}(s,a)=\frac{1}{n}\sum_{x=1}^{n}\bigg{(}\sum_{i=1}^{k-1}(\eta_{m}^{i}-\eta_{m-1}^{i})r_{i}^{(x)}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k}^{(x)},a_{n+k}^{(x)})+\eta_{m}^{k_{m}}\hat{D}_{m}(s_{n+k}^{(x)},a_{n+k}^{(x)})$ (29)
We will now ascertain the highest limit of the phased error. The cumulative errors of each D component, denoted as $TD(\Delta)$, is expressed as $\sum_{m=0}^{M}\Delta_{n}^{m}$, where $\Delta_{n}^{m}=\max_{s,a}\bigg{\{}\bigg{|}\hat{D}_{m}(s,a)-D_{m}(s,a)\bigg{|}\bigg{\}}$. Fundamental differences between SARSA and Q-learning regarding the error bound: In SARSA, the error is ascertained by the agent’s actual actions, indicating that the Q-values are modified based on the state-action pairings encountered by the agent during exploration. The inaccuracy in Q-learning arises from the greedy action, defined as the action that maximizes the Q-value for the subsequent state, based on the premise that the agent consistently behaves optimally.
@@
Inspired by Romoff et al.[8], assume that $\eta_{0}\leq\eta_{1}\leq\eta_{2}\leq.....\leq\eta_{m}=\eta$ and $K_{0}\leq K_{1}\leq...\leq K_{m}=K$, for any $10<\delta<$ 1, let $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$, with probability $1-\delta$,
$\sum_{m=0}^{M}\Delta_{n}^{m}\leq\epsilon\bigg{(}\frac{1-\eta^{k}}{1-\eta}\bigg{)}+\epsilon\bigg{(}\underbrace{\sum_{m=0}^{M-1}\frac{\eta_{m}^{k_{m+1}}-\eta^{k_{m}}_{m}}{1-\eta_{m}}\bigg{)}}_{\text{variance reduction}}+\underbrace{\sum_{m=0}^{M-1}\bigg{(}\eta^{k_{m}}_{m}-\eta^{k_{m+1}}_{m}\bigg{)}\sum_{q=0}^{m}\Delta_{n-1}^{q}}_{\text{bias introduction}}+\eta^{k}\sum_{m=0}^{M}\Delta_{n-1}^{m}$ (30)
@@
The author examines the comparison of the limits for phased TD($\lambda$) in theorem 3 with those for phased TD($\Delta$) in theorem 4. This comparison demonstrates that phased TD($\Delta$) facilitates variance reduction equivalent to $\epsilon\sum_{m=0}^{M-1}\frac{\eta_{m}^{k_{m+1}}-\eta^{k_{m}}_{m}}{1-\eta_{m}}\leq 0$ but introduces a potential bias quantified as $\sum_{m=0}^{M-1}\bigg{(}\eta^{k_{m}}_{m}-\eta^{k_{m+1}}_{m}\bigg{)}\sum_{q=0}^{m}\Delta_{n-1}^{q}\geq 0$. Utilizing phased TD($\Delta$) to decrease $k_{m}$ synchronizes updates more effectively with recent actions, hence lowering variance from high-discounted return components. Shortened $k_{m}$ values accrue bias more swiftly, as SARSA bootstraps from current policy estimations that may vary over phases. It is observed that when all $k_{m}$ values are identical, both algorithms yield the same upper bound.
@@
In SARSA, the return estimate encompasses subsequent states and actions dictated by the policy. The expected discounted return over T steps closely approximates the infinite-horizon discounted expected return after $\mathrm{T}$, where $\mathrm{T}\approx\frac{1}{1-\eta}$[31]. Consequently, we can effectively simplify $k_{m}$ for any $\eta_{m}$ such that $k_{m}\approx\frac{1}{1-\eta_{m}}$, thereby adhering to this principle. Therefore, with $\mathrm{T}$ samples, similar to TD($\Delta$)[8], SARSA employing TD($\Delta$) utilizes $k_{m}\approx\frac{1}{1-\eta_{m}}$ to establish a suitable balance between bias and variance over all time scales significantly less than $\mathrm{T}$. By establishing $k_{m}$ for each $\eta_{m}$, where $\eta_{m}^{\frac{1}{1-\eta_{m}}}\leq\frac{1}{e}$, the approach guarantees that the effective horizon and variance are maintained within limits without requiring extensive parameter optimization. Each increase in $D_{m}$ facilitates the doubling of the effective horizon, hence ensuring logarithmic scaling with the quantity of action-value functions. SARSA can utilize this to modify additional parameters in alignment with the specified time scales, enhancing the algorithm’s efficiency in calculating long-term rewards without added complexity.
6 Experimental Setup
We experimented with OpenAI Gym control tasks, and Fig.1 shows the relevant environments. The experiments were carried out on a single PC with 16 GB of memory, an Intel Core i7-2600 processor, and a GPU. The system ran 64-bit Ubuntu 22.04.5 LTS. We utilized Python 3.9 and the PyTorch package for development. We also used the OpenAI Gym toolkit, as mentioned by G. Brockman et al. [brockman2016openai]. Ten randomly selected seeds were used in each experiment, and the average outcomes are presented.
(a) FrozenLake-v1
(b) Taxi-v3
(c) Blackjack-v1
(d) CliffWalking-v0
Figure 1: OpenAI Gym control used for all experiments. In order from the left to right: Frozen Lake, Taxi, Blackjack and Cliff Walking.
6.1 Tabular
We utilize Q-value iteration to determine the optimal Q-value function. We employ an open-source version of Value Iteration (VI) derived from the code of Peter Henderson and Wei-Di Chang[value-iteration], guided by the algorithmic principles outlined in Bellman[22] and Sutton & Barto[29]. Although their methodology calculates the optimal value function, we modify it to determine the optimal Q-value function. Upon acquiring the optimal Q-values, we employ Temporal Difference (TD) and SARSA($\Delta$) to derive approximation Q-values via sampling. After 5000 steps, each learning iteration comes to an end. For every setup, we run 200 trials using different random seeds that influence the dynamics of the transition. We implement our algorithms using the code that Kostrikov provided[pytorchrl]. We compare our approach to k-step SARSA, as described in Sutton and Barto’s book, Chapter 7.2[29]. We evaluate SARSA($\Delta$) on multiple RL benchmarks, such as the Ring MDP[9] and deterministic & stochastic environments, to determine its effectiveness. A illustration of the Ring MDP is shown in Fig. 2, which shows a Markov process with five states arranged in a circle. There is a 0.95 chance of moving one state clockwise around the ring and 0.05 chance of staying in the present state at each step. While the other states on the ring offer a reward of 0, two adjacent states offer rewards of +1 and -1, respectively.
Figure 2: Five-state ring MDP as described by Kearns[9]
All experiments utilize the identical parameters established by Romoff et al.[8], which involve supplying a variable number of gamma values, commencing at 0 and incrementing according to $\eta_{m+1}=\frac{\eta_{m}+1}{2}$ until the requisite maximum $\eta_{m}$ is attained. As previously mentioned in Section 5.2, $K_{m}=\frac{1}{(1-\eta_{m})}$ for every m. The baseline consists of a singular estimator with $\eta=\eta_{m}$ and $k=k_{m}$. Fig. 3(a) depicts the efficacy of two methodologies: K-step SARSA ($\Delta$) and K-step SARSA TD, employing a discount factor ($\eta$ = 0.996), k = 64 (a parameter associated with the algorithm’s step count), and diverse values of the learning rate $\alpha$. The y-axis denotes the absolute error, whereas the x-axis signifies the parameter $\alpha$, ranging from 0.0 to 1.0. At $\alpha$=0.0, both methodologies commence at an equivalent error level (about 0.00230). As $\alpha$ increases, K-step SARSA ($\Delta$) initially diminishes its inaccuracy marginally before seeing a consistent rise. K-step SARSA TD, conversely, exhibits a consistent increase in error with minimal fluctuation. Overall course: Relative to K-step SARSA TD, the K-step SARSA ($\Delta$) technique has greater stability and consistently yields superior outcomes (i.e., reduced error) across all values of $\alpha$. Both approaches exhibit sensitivity to variations in $\alpha$, although K-step SARSA TD demonstrates greater sensitivity, evidenced by its more pronounced increase in error.
Fig. 3(b) demonstrates the effectiveness of the two methodologies, K-step SARSA ($\Delta$) and K-step SARSA TD, in relation to the number of k-steps. The y-axis indicates the absolute error, whereas the x-axis represents k-steps, which refers to the number of steps utilized in the algorithms. The range of k-steps extends from 0 to roughly 250. K-step SARSA ($\Delta$) commences with a reduced error at smaller k-steps, which subsequently increases with bigger k-steps. K-step SARSA TD begins with an error level comparable to K-step SARSA ($\Delta$) for low k-steps; however, its error escalates more steeply as k-steps grow. The overall pattern indicates that both techniques exhibit escalating inaccuracy as k-steps rise. K-step SARSA ($\Delta$) consistently surpasses K-step SARSA TD in all k-step values by exhibiting a reduced error rate. The swift error amplification in K-step SARSA TD indicates that the scheduling method employed in K-step SARSA ($\Delta$) mitigates error accumulation. Furthermore, in this specific MDP, the error escalates with k due to the bias-variance trade-off associated with k-step returns, aligning with the conclusions of Kearns & Singh[9, 8].
Comparison of Figs. 3(a) and 3(b): In the comparison of the two figures, K-step SARSA ($\Delta$) consistently surpasses K-step SARSA TD in both circumstances ($\alpha$ and k-steps). Both approaches exhibit sensitivity to elevated parameter values ($\alpha$ or k-steps). K-step SARSA ($\Delta$) exhibits enhanced stability and attains reduced errors as a result of its scheduling approach. Significantly, errors escalate more rapidly with K-step SARSA TD as k-steps grow in comparison to the increase in $\alpha$.
(a) Error at different learning rates
(b) Error at different k values
Figure 3: Fig.3(a) illustrates the outcomes for $\eta_{M}=0.996$ on the five-state Ring MDP. The error signifies the absolute discrepancy between the estimated and actual discounted Q-value function (precomputed via Q-value iteration), averaged across the complete learning trajectory of 5000 timesteps. Fig.3(b) illustrates the average absolute error for the optimal learning rate at each k-step return, extending beyond the effective planning horizon of $\eta_{M}$.
6.2 OpenAI Gym Environments
The following plots show the results of two algorithms, k-step SARSA($\Delta$) and k-step SARSA TD, on the FrozenLake-v1, Taxi-v3, CliffWalking-v0, and Blackjack environments from OpenAI Gym. The y-axis shows the average rewards obtained by the algorithms over episodes, and the x-axis shows the number of episodes needed to train the algorithms. Episodes range from 0 to 100,000 episodes. The shaded area surrounding each curve shows the standard error over 10 random seeds. The algorithms both use k= 32, which is the number of steps. The remaining hyperparameter settings are displayed in the table 1.
Table 1: Hyperparameter configurations and their corresponding values.
Hyperparameter Value
seed (10) np.random.randint(40, 45)
Learning Rate ($\alpha$) [0.1, 0.3, 0.5, 0.7, 1.0]
Discount Factor ($\eta$) [0.992, 0.95, 0.996, 0.99, 1.0]
k-Step [1, 4, 8, 16, 32, 128, 256]
Number of Episodes 100000
Max Epsilon 0.4
The red line denotes K-step SARSA ($\Delta$), while the blue line represents K-step SARSA TD in both Fig. 4(a) and Fig. 4(b). Fig. 4(a) illustrates the FrozenLake-v1 environment, whereas Fig. 4(b) depicts the Taxi-v3 environment. Fig. 4(a) demonstrates that k-step SARSA ($\Delta$) continually attains superior average rewards relative to k-SARSA TD as training advances. K-step SARSA ($\Delta$) demonstrates considerable enhancement in average rewards over the initial 20,000 episodes, signifying accelerated learning, but k-step SARSA TD exhibits lesser increase in average rewards in comparison to k-step SARSA ($\Delta$) during the same episodes. Following 40,000 episodes, the efficacy of k-step SARSA($\Delta$) stabilizes while consistently surpassing that of k-step SARSA TD, indicating enduring reliability. k-step SARSA regularly exhibits inferior performance relative to k-step SARSA ($\Delta$).
Fig. 4(b) illustrates that K-Step SARSA ($\Delta$) regularly achieves superior rewards (i.e., nearer to -200) during the training phase, signifying enhanced learning efficiency and overall performance. Fig. 4(b) also demonstrates that K-Step SARSA ($\Delta$) stabilizes in the initial phases, indicating robustness and the capacity to converge to an effective policy, whereas K-Step SARSA TD commences at considerably lower rewards (approximately -1800) and exhibits negligible improvement, remaining predominantly flat throughout the training process. The K-Step SARSA TD algorithm encounters difficulties in acquiring an effective policy in this setting compared to K-Step SARSA ($\Delta$). Experiment result show that K-Step SARSA ($\Delta$) significantly outperforms K-Step SARSA TD in the Taxi-v3 environment, achieving much higher average rewards and faster convergence.
(a) FrozenLake-v1
(b) Taxi-v3
Figure 4: Comparative analysis of K-step SARSA ($\Delta$) versus traditional K-step SARSA TD within the OpenAI Gym environments: FrozenLake-v1 and Taxi-v3. The shaded area illustrates the standard error calculated from 10 random seeds.
(a) CliffWalking-v0
(b) Blackjack-v1
Figure 5: Comparative analysis of K-step SARSA ($\Delta$) versus traditional K-step SARSA TD within the OpenAI Gym environments: CliffWalking-v0 and Blackjack-v1. The shaded area illustrates the standard error calculated from 10 random seeds.
The Fig. 5(a) and 5(b) illustrates the learning effectiveness of two variations of the K-Step SARSA algorithm, namely K-Step SARSA ($\Delta$) and K-Step SARSA TD, across 100,000 episodes in the CliffWalking-v0 and Blackjack-v1 environments. Both algorithms are assessed using k=32, a hyperparameter that specifies the number of steps in the updating process. The trend of average rewards in fig. 5(a) demonstrates that both algorithms show an increasing pattern in average rewards as the number of episodes rises, indicating improved policy performance with continued learning. Nevertheless, K-Step SARSA ($\Delta$) often outperforms K-Step SARSA TD, achieving higher average rewards over the episodes.
The red line in fig. 5(a) denoting K-Step SARSA ($\Delta$) demonstrates a considerable enhancement in the initial episodes (about 20,000 episodes), indicating a swifter convergence to optimal policies relative to K-Step SARSA TD. In the 100,000th episode, K-Step SARSA ($\Delta$) attains a reward near -1050, whereas K-Step SARSA TD obtains a slightly lower payout of approximately -1120. This demonstrates the enhanced efficacy of K-Step SARSA ($\Delta$) in this task.
The average reward of approximately +0.15 is dependably achieved across all episodes by the red curve in fig. 5(b) that represents K-Step SARSA ($\Delta$). The blue curve in fig. 5(b), which represents K-Step SARSA TD, suggests stability; however, it presents an average reward that is significantly lower, at approximately -0.2, across all episodes. K-Step SARSA ($\Delta$) markedly surpasses K-Step SARSA TD regarding average rewards. The favorable average reward of K-Step SARSA ($\Delta$) demonstrates its superior suitability for policy optimization in the Blackjack-v1 environment, whereas the negative reward for K-Step SARSA TD signifies inferior performance.
Finally, when comparing the results of all four environments, the findings demonstrate the constant advantage of K-Step SARSA ($\Delta$) over K-Step SARSA TD across all evaluated environments. This superiority is seen through elevated average rewards, accelerated convergence, and enhanced stability in learning outcomes. Environments characterized by deterministic state transitions, such as CliffWalking-v0 and FrozenLake-v1, demonstrate the clear superiority of K-Step SARSA ($\Delta$), but intricate stochastic environments like Blackjack-v1 and Taxi-v3 further highlight its resilience. The performance graph of deterministic environments, such as CliffWalking-v0 and FrozenLake-v1, resembles an exponential growth plateau curve due to predictable state transitions. In contrast, the performance graph of complex stochastic environments, like Blackjack-v1 and Taxi-v3, appears as a flat line because of the uncertainty and variability in state transitions.
@@ 7 Conclusion
With the help of TD($\Delta$), SARSA($\Delta$) improves SARSA by integrating multiple discount factors, leading to better stability and performance. With this approach, we can get a more precise breakdown of action-values while still enjoying the benefits of on-policy learning. In addition to the deterministic and stochastic OpenAI Gym environments, we also conducted trials in a tabular synthetic environment, specifically the five-star ring MDP. In all investigated environments, our experimental results demonstrate that SARSA($\Delta$), our suggested approach, outperforms k-step SARSA TD in terms of greater average rewards, faster convergence, and more stable learning trajectories. The results of this work advance our understanding of multi-step RL algorithms and how they are applied to various sequential decision-making tasks. This information can aid in the creation of robust algorithms for unpredictable, real-world environments in the future. In essence, the Q-Learning version of TD($\Delta$) extends this technique. We decided to take care of it later.